A balloon containing 0.56 mol of helium has a volume of 10.0 L. What volume will the balloon occupy if helium is lost and only 0.21 mol remain?
1
Expert's answer
2013-02-27T06:00:27-0500
According to the ideal gas law
pV1 = n1RT - at the beginning

pV2 = n2RT - after part of helium is lost

p (pressure), T(temperature), R (universal gas constant) - don't change
V1, V2 - volumes
n1, n2 - moles

(pV1/pV2) = (n1RT)/(n2RT)

V1/V2 = n1/n2

After helium is lost the volume is

V2 = (V1n2)/n1

V2 = (10.0 x 0.21)/ 0.56 = 3.75 L

Answer: V2 = 3.75 L